OpenStudy (marcoreus11):

I can across this problem on my sat prep book can someone teach me how to do it? I don't want the answer lol if I wanted I would of went to the back of the book:D

OpenStudy (marcoreus11):

Here

OpenStudy (marcoreus11):

@Seratul

zepdrix (zepdrix):

Can you check the back of the book real quick or no? :) I don't want to explain this if I did it incorrectly. Is the answer c=3?

OpenStudy (marcoreus11):

yes

OpenStudy (marcoreus11):

it is c=3

zepdrix (zepdrix):

Since the polynomial is divisible by y-2, it means we can divide y-2 out of the polynomial,\[\large\rm y^5-2y^4-cxy+6x\quad=\quad (y-2)(stuff)\]And it will leave us with some other polynomial stuff, but no remainder. What happens if we plug y=2 into this new thing?\[\large\rm (y-2)(stuff)=(2-2)(stuff)\]The whole thing ends up giving us zero, right? That's what our remainder theorem tells us. Our function evaluated at y=2 will give us zero because the expression is divisible by y-2.

zepdrix (zepdrix):

\[\large\rm y^5-2y^4-cxy+6x\]So when we plug in y=2, we can set this expression equal to zero,\[\large\rm 2^5-2(2)^4-cx(2)+6x=0\]

OpenStudy (marcoreus11):

okay so basically for these problems we set the problem to 0 and plug in the numbers for y?

zepdrix (zepdrix):

For this specific type of problem, where we have divisibility, yes it looks like that will always be the case! :)

OpenStudy (marcoreus11):

okay so if its y+2 it would be y=-2?

zepdrix (zepdrix):

Yes

OpenStudy (marcoreus11):

oh okay so any type of problems that says polynomial divisible by we just make it opposite and plug it into the letter and equal to zero.

OpenStudy (marcoreus11):

just wanted to make sure and thank you

zepdrix (zepdrix):

Yes. It's helpful if you understand `why` this is true though. That's what I was trying to do with the (y-2)(stuff) explanation. But as long as you have some type of shortcut memorized, that will work out nicely. :) np

OpenStudy (marcoreus11):

i was worrying so much about it a shortcut lol we didn't solve it @zepdrix i tried to simplify it but i ended up something else then 3

OpenStudy (marcoreus11):

if u could help me with that, i would appreciate it thx

zepdrix (zepdrix):

\[\large\rm 2^5-2(2)^4-cx(2)+6x=0\]2*2^4 is the same as 2^5, right? So the first two terms subtract off,\[\large\rm -cx(2)+6x=0\]

OpenStudy (marcoreus11):

yes

zepdrix (zepdrix):

Let's rewrite our equation like this:\[\large\rm 6x-2cx=0\]From here we should try to factor. I guess we can pull an x out of each term, and maybe a 2 also, ya?

OpenStudy (marcoreus11):

oh so u canceled out the x

OpenStudy (marcoreus11):